If the population of a town increases every year by 2% ofa total population at the begining of that year, in how many year will the total increase of population be 40%?
Answer by Boreal(15235)

original is Po, final is P
P=Po{1+.02}^n, the interest formula.
P/Po=1.40
1.40={1.02}^n
ln of both sides
ln(1.40)=n{ln (1.02))
divide by ln 1.02
ln(1.40)/ln(1.02)=16.99 years
17 years.
1.02^17=1.400